# Dynamical systems approach to Cold and Warm Inflation within slow-roll and beyond

### v1

@@ 5.2 CRWI with $\beta<0$ and constant $Q$
$P_{1}$ $(0,0)$infty $\infty$ $-\frac{3Q}{2}-1$ Unstable
$P_{2}$ $(\rm{Any},1)$ 00 $2$ $-1$ Unstable
@@
Figure 11: The phase space corresponding to CRWI with $\beta<0$. The color scheme highlights the different dynamical regions: pink ($\epsilon_{1}\geq 1$), yellow region ($0\leq\epsilon_{1}\leq 0.1$), green region ($10.1<\epsilon_{1}<$ 1), blue (thermally stable region), and violet (overlap of red and blue regions).
The phase space for the constant-roll warm inflation (CRWI) model is illustrated in Fig. [11] for various parameter choices of $(Q,\tilde{\beta})$ within the $(\bar{x},\bar{z})$ space. The phase space is depicted using distinct color codes: the pink region corresponds to $\epsilon_{1}\geq 1$, the yellow region corresponds to $0\leq\epsilon_{1}\leq 0.1$, the green region corresponds to $10.1<\epsilon_{1}<$ 1 and finally the blue region specifies the set of points in the phase space where we have thermal stability . Where the blue and red regions overlap, the color appears violet. Additionally, the white region represents areas where the slow-roll parameter is negative ($\epsilon_{1}<0$) and hence these regions are practically ruled out from any meaningful dynamical analysis. Since $\bar{z}=1$ corresponds to a line of unstable fixed points, Fig. [11] illustrates that trajectories diverge away from it. Additionally, the figure shows that all trajectories move toward $V=0$, i.e., the $\bar{x}=0$ line.
@@ 6 Conclusion
In this work we have tried to formulated a methodical approach to study the qualitative as well as the quantitative features of the phase space of the standard inflationary models and their constant-roll variants. To set the stage we began our investigation with the well-known cold inflationary scenario, considering a quadratic potential for the scalar field. Utilizing a Poincaré transformation, we compactified the system and identified that the only fixed point corresponds to the coordinates $(\phi,\dot{\phi})=(0,0)$. Notably, even if the system starts in a non-inflationary region, it transitions into the inflationary phase, characterized by $\epsilon_{1}\ll 1$, and eventually exits when $\epsilon_{1}\sim 1$. These results are consistent with findings reported in the literature. We have focused on a morphological study of the qualitative features of phase space in various cases of inflationary dynamics where all kinds of phase trajectories are present. We have not purely concentrated on trajectories which follow slow-roll inflation although in all cases we have shown that some slow-roll trajectories do exist for some parameters of the theory and for some class of initial conditions.
In our formulation of phase space dynamics of inflationary systems we have introduced the concept of compactified phases space to identify all the possible fixed points of the system and this method also gives a qualitative feature of the global behavior of the phase space. This construction turns out to be important in inflationary dynamics as inflation occurs for a very, very brief time period whereas the variables which describes the system ae defined in much longer timescale. In such a case one has to be sensitive about the presence of stable fixed points inside the inflationary regime, if such fixed points do exist then they can be dangerous for inflation. Consequently, for a proper inflationary dynamics one requires the stable fixed points to be in the asymptotic past or in the asymptotic future of the inflationary phase. Only unstable fixed points are allowed to exist in the regions of phase space where the slow-roll parameter $\epsilon_{1}<1$. The compactification scheme helps us to identify the fixed points at the asymptotic regions.
We have addressed the issue of singularities in the autonomous equations for the inflationary system. It has been shown that the singularities arising in the autonomous set of equations can be regulated by introducing a new time variable which is a monotonic increasing function of the old time variable. The regulation technique is mostly used in cosmological models of dark energy, for inflationary dynamics we have successfully used this powerful technique for the first time. We have shown that the duration of separation of events in the phase space do depend upon the choice of the time parameter. Using an inverse transformation one can go back to the standard time coordinates after the calculations are done. Using this technique we have found out the phase space dynamics in WI and CRWI.
@@
Compared to standard cold inflation and warm inflation models, the constant-roll variants of these models are relatively constrained. The constraint comes from the constant-roll condition and is dynamic in nature. This dynamic constraint has appreciable effect on the dynamics as it drastically affects the dimension as well as the morphology of the phase space. The constant-roll condition directly affects the issue of graceful exit from inflation. Moreover the constant-roll condition transforms the dynamical system of CI into an algebraic system where one can figure out the phase space dynamics without explicitly solving any set of autonomous equations. Due to the constant-roll condition, the actual phase space of CRCI is one dimensional. To make the results more tractable we have plotted a recreated two dimensional stream-plot where the stramlines have phase space coordinateds in the case of CRCI. Compared to CRCI, CRWI has a genuine two dimensional phase space and CRWI can be tackled both algebraically as well as with the system of autonomous differential equations. In the present paper we have chosen the system of autonomous equations to study the dynamics of CRWI.
In the paper we have also established a standard method of studying stability analysis in the case of warm inflation. We present the result of warm inflation as it is a precursor of CRWI. We have dealt the case of warm inflation with utmost sensitivity as because the dynamics of the warm inflationary process is a delicate one as many conditions have to simultaneously satisfied during WI. These various conditions were discussed in the introduction. For the first time we have proposed a very general dynamical systems approach which can take care of all the subconditions in WI and produce the general structure of the phase space in WI. Using the knowledge gained in dealing the dynamical systems approach in WI we have modelled the produced the phase space structure of CRWI. Our work shows that the phase space of CRWI dynamics is nontrivial as due to various conditions applied on the system a part of the phase space remains dynamically unavailable.
@@
Inflationary phenomenology is in general aimed at producing 60-70 $e$-folds of inflation and thereby producing cosmological perturbations so that structures can be formed in the later part of cosmological evolution. More emphasize is given to the calculation of power spectrum which is generated from various inflationary models and comparing those results with observational data. The dynamical stability and attractor nature of the background model in inflation is studied rarely. This study is very important as because the inevitability of cosmological inflation greatly depends upon the attractor nature of the inflationary processes in the phase space. In this paper we have introduced a thorough and methodical way to systematically study the dynamical structure and stability of the background inflationary models. We hope in the near future we will also be able to propose a methodical way to study the growth of inflationary perturbations.
@@
Acknowledgement: The authors acknowledge professor Suratna Das’ comments regarding various points discussed in this paper.

### v3

@@ 3.2 CRWI with $\beta<0$ and constant $Q$
$P_{1}$ $(0,0)$ $\infty$ $\infty$ $\frac{3Q}{2}-1$ Unstable
$P_{2}$ $(\rm{Any},1)$ 0 22 1-1 Unstable
@@
Figure 8: The phase space corresponding to CRWI with $\beta<0$. The color scheme highlights the different dynamical regions: pink ($\epsilon_{1}\geq 1$), yellow region ($0\leq\epsilon_{1}\leq 0.1$), green region ($0.1<\epsilon_{1}<1$), blue (thermally stable region), and violet (overlap of red and blue regions).
[…]
@@
4 Warm ultraslow-roll (USR) inflation and its stability
In warm inflation, an ultraslow-roll (USR) phase refers to a period during which the inflaton rolls on an almost flat section of the potential, but unlike standard slow-roll, the acceleration term $\ddot{\phi}$ is not negligible. In fact, it plays a crucial role in the dynamics. The inflaton equation of motion in warm inflation takes the form
$\ddot{\phi}+(3H+\Upsilon)\dot{\phi}+V_{,\phi}=0,$ (4.1)
where $\Upsilon$ is the dissipation coefficient. During USR, the potential gradient $V_{,\phi}\equiv dV/d\phi$ is much smaller than the friction and the acceleration term, so the field evolution is driven mainly by the damping terms and its own acceleration, i.e., $\ddot{\phi}+(3H+\Upsilon)\dot{\phi}=0$. Compared to the cold inflation case, the presence of $\Upsilon$ adds an extra source of friction, which can affect both the duration and stability of the USR phase. Depending on the strength of dissipation, the system may linger in USR longer, or exit it more quickly. Understanding USR in warm inflation is especially interesting since it can affect the generation of curvature perturbations, and may have implications for features like enhanced power spectra or even primordial black holes[36, 82, 64].
4.1 Dynamical analysis of warm USR inflation in the potential $V(\phi)=V_{0}+M^{3}\phi$
In Ref. [20], the authors explored the scope of warm USR inflation with a linear potential of the form:
$V(\phi)=V_{0}+M^{3}\phi\ ,$ (4.2)
where $V_{0}$ is a density constant, and $M$ is a mass scale. Additionally, it has been assumed that $\kappa V_{0}\gg M^{3}$ consequently $V_{0}\gg M^{3}\phi$. Under this condition, the dynamics of USR has been studied.
In this section, we explore the dynamics of the present scenario using the dynamical systems stability framework. The presence of higher-derivative terms in the current model may lead to novel dynamical features that have not been addressed in previous studies. Following the same approach as before, we define the dimensionless variables:
$x\equiv\sqrt{\frac{V(\phi)}{3V_{0}}},$ $y\equiv{\frac{\dot{\phi}}{\sqrt{6V_{0}}}},$ $z\equiv\frac{H}{\kappa\sqrt{V_{0}}}\ .$ (4.3)
Ref. [20] demonstrated that achieving ultra-slow-roll (USR) inflation within the warm inflation (WI) framework requires the dissipation coefficient $\Upsilon$ to depend, at minimum, on the inflaton field $\phi$, i.e., $\Upsilon=\Upsilon(\phi,T)$. In particular, Ref. [20] considered the form $\Upsilon(\phi,T)=C_{\Upsilon}\frac{T^{3}}{\phi^{2}}$, where $C_{\Upsilon}$ is a dimensionless constant. The equations governing the WI in this scenario are:
$\ddot{\phi}+3H\dot{\phi}+C_{\Upsilon}\frac{T^{3}}{\phi^{2}}\dot{\phi}+M^{3}$ $=0\ ,$ (4.4)
$\dot{\rho}_{r}+4H\rho_{r}$ $=C_{\Upsilon}\frac{T^{3}}{\phi^{2}}\dot{\phi}^{2}\ .$ (4.5)
Using the Friedmann equation (Eq. 3.7), we can express the radiation energy density in terms of our dimensionless variables:
$\frac{\rho_{r}}{3V_{0}}=z^{2}-y^{2}-x^{2}.$ (4.6)
Now we construct the autonomous system of equations in terms of the dimensionless dynamical variables using the field equation as:
$x^{\prime}=\frac{M^{3}\sqrt{6}}{\kappa V_{0}}\frac{y^{2}}{x}\ ,$ (4.7)
$y^{\prime}=-\left(3yz+C_{\Upsilon}C_{R}^{3/4}27^{1/4}M^{6}\frac{(z^{2}-y^{2}-x^{2})^{3/4}}{\kappa V_{0}^{7/4}(3x^{2}-1)}+\frac{M^{3}}{\kappa V_{0}}\right)\ ,$ (4.8)
$z^{\prime}=2x^{2}-y^{2}-2z^{2}\ .$ (4.9)
where $()^{\prime}\equiv\frac{1}{\kappa\sqrt{V_{0}}}\frac{d}{dt}$ and we define $t^{\prime}=\kappa\sqrt{V_{0}}t$. For this scenario, the phase space becomes 3-dimensional, however,the system does not have any critical points. Hence, we will conduct our further analysis numerically. Before we do, at first we express few important quantities in terms of dynamical variables. For instance, the Hubble slow roll parameter $\epsilon_{1}\equiv-\frac{\dot{H}}{H^{2}}$ can be expressed in terms of dynamical variables as:
$\epsilon_{1}$ $=\frac{2x^{2}-y^{2}-2z^{2}}{z^{2}}\ ,$ (4.10)
and similarly, the second slow roll parameter can be expressed as
$\epsilon_{2}\equiv\frac{\dot{\epsilon_{1}}}{H\epsilon_{1}}=2\epsilon_{1}+\frac{4xx^{\prime}-2yy^{\prime}-4zz^{\prime}}{z(2x^{2}-y^{2}-2z^{2})}\ ,$ (4.11)
The ultra-slow-roll (USR) phase is characterized by the conditions $\epsilon_{1}<1$ but $|\epsilon_{2}|>1$. In case of WI, the system must satisfy $|\frac{1}{H}\frac{\dot{\rho}_{r}}{\rho_{r}}|<1$. Once the USR phase ends after a few e-folds, the inflaton field transition to the slow-roll (SR) phase, marked by $\epsilon_{1}<1$ and $|\epsilon_{2}|<1$. We use a plot showing the variation of $y\text{--}z$ in Fig. [10], because the full three-dimensional trajectory does not provide clear information about whether the system converges to any particular region in phase space. In contrast, the 2D parametric plot offers more insight into the USR phase and other dynamical regimes. However, the curves in this plot do not appear to converge to any particular region in the phase space, indicating that the USR phase is not a stable attractor. In the parametric plot, we show three distinct trajectories corresponding to three different sets of initial conditions. These are represented by the orange, blue, and green curves, each following noticeably different evolutionary paths during the USR regime. The parameters used are $V_{0}=(10^{-4})^{4}M^{4}_{\mathrm{pl}}$, $M=2.5\times 10^{-7}M_{\mathrm{pl}}$, $C_{\Upsilon}=10^{5}$, and $g_{*}=106.75$. The orange trajectory starts from $x(0)=5.77365\times 10^{-1}$, $y(0)=1.4506\times 10^{-3}$, $z(0)=5.77366\times 10^{-1}$; the blue trajectory begins at $x(0)=5.781\times 10^{-1}$, $y(0)=1\times 10^{-3}$, $z(0)=5.782\times 10^{-1}$; and the green trajectory originates from $x(0)=5.9\times 10^{-1}$, $y(0)=4\times 10^{-3}$, $z(0)=6.01\times 10^{-1}$. As time evolves, all three trajectories moves leftward in the $y\text{--}z$ plane, yet none of them converge to a common point, reinforcing the interpretation that the USR phase does not correspond to a stable solution. Moreover, we observe only portions of these trajectories in the $y\text{--}z$ plane because the system of differential equations becomes stiff after a certain time, making numerical integration impossible beyond that point.
Figure 10: Phase-space dynamics in the $y$-$z$ plane showing orange, blue, and green trajectories corresponding to three different initial conditions. The colored paths reveals the non-attractor behavior intrinsic to USR inflation.
To study the initial condition of the system and the fate of warm USR it is more suitable if we follow the approach of Ref. [82]. In warm inflationary scenario we can generalize the treatment in [82] and define the dimensionless parameters:
$f_{\phi}=-\frac{\ddot{\phi}}{3H\dot{\phi}+\Upsilon\dot{\phi}},$ $f_{\rho}=\frac{\dot{\rho}}{4H\rho}\ .$ (4.12)
In Warm Ultra-Slow-Roll (USR) regime, these parameters exhibit $f_{\phi}\simeq 1$,and $|f_{\rho}|<1$, reflecting production of radiation energy maintaining thermal equilibrium. These parameters distinguishes between different inflationary regimes: (i) slow-roll (SR) regime: $|f_{\phi}|\ll 1$ and (ii) ultra-slow-roll (USR) regime: $|f_{\phi}-1|\ll 1$, while maintaining $|f_{\rho}|<1$ through out the evolution in both cases. We linearise $f$ about $f=1$ by parameterising
$f_{\phi}=1-\delta\ ,$ (4.13)
where $|\delta|\ll 1$. As $|\delta|$ increases (decreases) with time the USR becomes unstable (stable). As the predefined variables can take any arbitrary values, now we compactify those using the Poincaré transformation
$x\to\frac{x}{\sqrt{1+x^{2}}},$ $y\to\frac{y}{\sqrt{1+y^{2}}},$ $z\to\frac{z}{\sqrt{1+z^{2}}},$ (4.14)
resulting in $(x,y)\in[-1,1]$ and $z\in[0,1]$ range. Compactifying the variables helps us to choose the initial conditions as now we can in principle choose any allowable points in the compactified phase space as initial conditions. If we have not compactified the phase space, points at infinity or points far away from the origin may have been left out.
For the given potential in Eq. (4.2), the system cannot start in the SR regime where $\delta\sim 1$. Because, if we choose the initial conditions corresponding to such cases, the system gets stuck in the SR regime which is the attractor phase and unable to render the USR phase. This is our main prediction about the initial condition in warm USR inflation. In a certain sense warm USR is very different from other models of inflation as here one cannot have a transient warm USR phase terminating in a SR phase unless one starts with initial conditions which do satisfy USR conditions, hence we may infer that such a transcient phase lacks an attractor solution as general initial conditions are not attracted towards warm USR. Hence, we choose the initial conditions corresponding to the USR phase i.e., $\delta\ll 1$. Now we plot $f_{\phi}$ for the transformed variables in Fig. [11]. The results indicate optimal behavior when the initial values are given as $x_{0}\approx 0.51$, with $z_{0}\gtrsim x_{0}$ and $y_{0}\ll|x_{0}|$. As we increase $x$ beyond this point, the system spends progressively less time in the USR phase, marked by $f_{\phi}\sim 1$ and transit to $f_{\phi}\to 0$, SR phase. For $x_{0}<0.51$, $f_{\rho}<-1$, which violates the energy conditions.
Figure 11: Evolution of the dissipation coefficient $f_{\phi},f_{\rho}$ for a linear potential $V(\phi)=V_{0}+M^{3}\phi$. The limit $f_{\phi}\to 0$ signals the onset of slow-roll inflation. Also $f_{\rho}>-1$ indicates that the system in those cases evolves maintaining thermal equilibrium, while for the orange line, which corresponds to $x_{0}<0.51$, we have $f_{\rho}<-1$, which violates the condition $|f_{\rho}|<1$ necessary for thermalization.
To visualize the impact of initial conditions, we have plotted several distinct curves in the 2D plots in Fig. [11], each corresponding to a different set of initial values of $(x_{0},y_{0},z_{0})$, and each assigned a unique color. The initial conditions span the ranges $x_{0}\in[0.4,0.75]$, $y_{0}\in[-10^{-3},-10^{-2}]$, and $z_{0}\in[0.4,0.8]$. The brown, magenta, and red curves represent variations around a fiducial trajectory with small deviations in all three variables. The blue and green curves correspond to larger values of $x_{0}$, with $y_{0}$ and $z_{0}$ held near their fiducial values. The purple and cyan curves explore significant deviations in $z_{0}$, while keeping $x_{0}$ and $y_{0}$ fixed. The orange curve illustrates the effect of decreasing $x_{0}$, with $y_{0}$ and $z_{0}$ again held fixed. This color scheme allows us to track how variations in each parameter influence the evolution of $f_{\phi}$, particularly in identifying conditions that either extend the USR phase or lead to an earlier transition into the slow-roll regime. We observe that in all cases, the system eventually transitions toward the slow-roll attractor, indicating that the USR phase is a transient phase.
Previous authors [81, 26] have noted that cold USR phase is a non-attractor phase. In this work we have shown that even warm USR inflation is not the standard attractor solution as there may be initial states in the phase space (which do not obey the USR condition) which may not ultimately stabilize to a slow-roll phase after the USR phase. One may say the warm USR inflation is a particular form of non-attractor solution with more subtlety. The subtlety arises because one has to keep in account the thermalization condition. In this section it is specifically shown which kind of initial conditions can give rise to a warm USR phase followed by a slow-roll phase maintaining the thermalization condition. If we deviate from the thermalization condition the stability analysis will get modified. Our work shows the way in which the attractor solution fails in the present case and opines about the stability of warm USR phase of inflation.
@@ 5 Conclusion
In this work we have tried to formulated a methodical approach to study the qualitative as well as the quantitative features of the phase space of the nonstandard inflationary models. We have studied non-slow-roll inflationary models where one of the models do not have any slow-roll limit. Our primary emphasize has been on the possible initial conditions which can give rise to various non-slow-roll inflationary phases.
Our work presents some new results. Some of these new results are general and some are particular. In this paper we have for the first time shown that there can be a particular potential, as discussed in section 2, in which one can have CRCI. This result is new as previously the authors in Ref. [70] the authors opined that CRCI cannot occur in the aforementioned potential. We have specified the nature of initial conditions in CRCI for various values of the CRCI parameter $\beta$. Next we have for the first time discussed the phase space description of CRWI.
@@
Compared to standard cold inflation and warm inflation models, the constant-roll variants of these models are relatively constrained. The constraint comes from the constant-roll condition and is dynamic in nature. This dynamic constraint has appreciable effect on the dynamics as it drastically affects the dimension as well as the morphology of the phase space. The constant-roll condition directly affects the issue of graceful exit from inflation. Moreover the constant-roll condition transforms the dynamical system of CI into an algebraic system where one can figure out the phase space dynamics without explicitly solving any set of autonomous equations. Due to the constant-roll condition, the actual phase space of CRCI is one dimensional. To make the results more tractable we have plotted a recreated two dimensional stream-plot where the streamlines have phase space coordinateds in the case of CRCI. Compared to CRCI, CRWI has a genuine two dimensional phase space and CRWI can be tackled both algebraically as well as with the system of autonomous differential equations. In the present paper we have chosen the system of autonomous equations to study the dynamics of CRWI.
In the last section we have extended our approach to the case of warm USR inflation. In the USR variation of WI, we have shown that the system does not admit general initial conditions. Only those initial conditions are allowed here, for which we alrady have the USR condition fulfilled. Only these initial states can ultimately have a transient USR inflation phase which terminates in a SR inflation phase. From these findings it can be seen that entering an USR phase in WI is in general difficult and not this phase requires very particular kind of initial conditions. These results are obtained using a particular form of the inflaton potential, but we expect that the results are fairly general and will hold true for other kind of inflaton potentials giving rise to a transient USR phase.
@@
Inflationary phenomenology is in general aimed at producing 60-70 $e$-folds of inflation and thereby producing cosmological perturbations so that structures can be formed in the later part of cosmological evolution. More emphasize is given to the calculation of power spectrum which is generated from various inflationary models and comparing those results with observational data. The dynamical stability and attractor nature of the background model in inflation is studied rarely. This study is very important as because the inevitability of cosmological inflation greatly depends upon the attractor nature of the inflationary processes in the phase space. In this paper we have introduced a thorough and methodical way to systematically study the dynamical structure and stability of the background inflationary models for non-slow-roll models of inflation. We hope in the near future we will also be able to propose a methodical way to study the growth of inflationary perturbations.
Acknowledgement
S.H. acknowledges the support of National Natural Science Foundation of China under Grants No. W2433018 and No. 11675143, and the National Key Reserach and development Program of China under Grant No. 2020YFC2201503.
@@ Acknowledgement
The authors acknowledge professor Suratna Das’ comments regarding various points discussed in this paper.